OpenStudy (anonymous):

use long division. x^2+x+1/x^3-2x^2+5

OpenStudy (campbell_st):

do you have the problem written correctly.... as all you have is the remainder

OpenStudy (anonymous):

wait sorry its the other way around

OpenStudy (campbell_st):

\[\frac{x^3 -2x^2 + 5}{x^2 + x + 1}\]

OpenStudy (campbell_st):

 then you have the quotient \[Q(x) = x - 3\] the remainder \[R(x) = 2x + 8\]